Operations Report – January 23rd

Crew 203 Operations Report 23-JAN-2019

SOL: 11

Name of person filing report: Freddy Castaneda

Non-nominal systems: Nothing new to report.

Notes on non-nominal systems: None

Generator (hours run): Unknown; Turned on last night (22Jan2019) at 16:50; Stopped running during the night, no hour reported on the generator. It was turned on this morning for about 15 minutes (23Jan2019) from 10:10 to 10:25; Turned on tonight (23Jan2019) about 16:30

Solar SOC – Turned on (22Jan2019) 83%; Turned off (23Jan2019) 60%; Turned on (23Jan2019) 84%

Diesel Reading – Under 1/2 tank.

Propane Reading – Refilled, 80%

Ethanol Free Gasoline – Nothing to report.

Water (auxiliary tank) – Not in use.

Water (static tank) – About 30%; 160 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 43%; 130 gallons

Water (loft) – At level marker 10

Static to Loft Pump used – Yes; At 14:30 to refill the tank

Water Meter: 01404157

Toilet tank emptied: Yes.

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Used for EVA to North Pinto Hills.

Hours: 70.3

Beginning charge: 100%

Ending charge: 64%

Currently charging: Yes

Opportunity rover used: Not used.

Hours: 46.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Used for EVA to North Pinto Hills.

Hours: 72.2

Beginning charge: 100%

Ending charge: 89%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly.

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Not used.

CrewCar used and why, where? Off-campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios:

Suit #1: All electrical connections verified, working properly.

Suit #2: All electrical connections verified, working properly.

Suit #3: Negative electrical terminal to battery replaced, All electrical connections verified, fuse used to enable suit #9, the battery doesn’t make an entire cycle, charges and discharges completely in less than one minute. Battery retired for exchange.

Suit #4: Loose terminal screw on the switch, all electrical connections verified, working properly.

Suit #5: All electrical connections verified, working properly.

Suit #6: Fan locked, voltage on battery was down and battery discharges quickly, the second fan move slowly. Issue fixed, all electrical connections verified, working properly.

Suit #7: All electrical connections verified, working properly.

Suit #8: Negative electrical terminal on battery replaced, electrical fail on fan, issues fixed, all electrical connections verified, working properly.

Suit #9: Enable with fuse of suit #3 and battery of suit #11, all electrical connections verified, working properly.

Suit #10: Battery replaced with an old one taken from the RAM, all electrical connections verified, working properly. Pending for battery functionality on the field.

Suit #11: Battery replaced with an old one taken from the RAM, all electrical connections verified, working properly. Pending for battery functionality on the field.

Summary of Hab operations: 8 toilet paper rolls received.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support: Nothing to report.